OpenStudy (anonymous):

x^(4/5)(x-10) has two critical numbers where A

OpenStudy (anonymous):

What're critical numbers, when the whole expression gets to Zero?

OpenStudy (anonymous):

a critical point of a function occurs when the derivate is 0 or the function is not differentiable (when the derivate is undefined).

OpenStudy (anonymous):

Hmm \[f(x)=x^{4/5+1}-10x^{4/5}\]\[f'(x) = \frac 9 5x^{4/5} - 10x^{-1/5}\]\[f'(x) = 0\implies \frac 9 5x^{4/5} - 10x^{-1/5}=0\]

OpenStudy (anonymous):

Sorry, \(f'(x)= \frac 9 5x^{4/5} - 10\times \frac{4}{5}x^{-1}{5}\)

OpenStudy (anonymous):

I hope you can solve it now

OpenStudy (anonymous):

Typo :/ ...Sorry \[f'(x)= \frac 9 5x^{4/5} - 10\times \frac{4}{5}x^{-1/5}\]BTW at x = 0, the function is undefined.